A long distance phone call costs $0.96 for the first minute and $0.72 for each additional minute or portion thereof. Write an inequality representing the number of minutes a person could talk without exceeding $6.
Answer by 13114(5)

Let the number of minutes he could talk be 'x'.
For the first minute the cost is $0.96.
So, for the remaining 'x - 1' minutes the cost is $0.72.
So, the total cost is $0.72(x - 1) + 0.96.
The amount should not exceed $6.
So, the inequality is $0.72(x - 1) + 0.96 < = $6
That is, 0.72x + 0.24 < = $6
That is 72x + 24 < = 600.